Evaluate the logarithmic equation for three values of x that are greater than 1, three values of x that are between 0 and 1, and at x=1. Use the resulting ordered pairs to plot the graph. State the equation of the line asymptotic to the graph (if any).
y = log1/7 (x+2)
This is what I have so far, indeterminite form which can't be graphed, or can it?
y=(log^1)/(7(x+2)
x+2<=0
x<=-2
x>-2 (-2,I)
x=
I
L[x:I,(log^1)/(7(x+2)]
The value of L[x:I,(log^1)/(7(x+2)] is (log^1)/(7(I+2).
(log^1)/(7(I+2)
I
y=(log^1)/(7(I+2)
I
L[x:-I,(log^1)/(7(x+2))]
L[x:-I,(log^1)/(7(x+2)] is (log^1)/(7(Indeterminate Form).
Indeterminite form

Answer by stanbon(75887)

Evaluate the logarithmic equation for three values of x that are greater than 1, three values of x that are between 0 and 1, and at x=1. Use the resulting ordered pairs to plot the graph. State the equation of the line asymptotic to the graph (if any).
y = log1/7 (x+2)
---------------------------------
If you want to graph it use y = [log(x+2)]/[log(1/7)].
-----------------------------------
x=1 ; y =-0.5646
x = 1/3; y= -0.4354
x = 1/2; y = -0.4709
x = 3/4; y = -0.5199
x = 2; y = -0.7124
x = 3; y = -0.8271
x = 4; y = -0.9208
=========================
Cheers,
Stan H.